40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5205     Entry time: Fri Aug 12 11:07:50 2011
Author: Nicole 
Type: Summary 
Category: SUS 
Subject: More TT Shaking Completed This Morning 

This morning (about 10am to 11am), I have collected additional transfer function measurements for the T.T. suspension. I have finished taking my measurements. The SR785 has been returned to its place next the the seismometer racks.

 The data has been backed up onto the cit40m computer

ELOG V3.1.3-